Referring to the information below, what is the probability that he will receive anywhere from two to six calls inclusive?
The probabilities that a doctor's answering service will receive 0, 1, 2, 3, 4, 5, 6, or 7 calls for him during lunch hour are, respectively, 0.001, 0.006, 0.022, 0.052, 0.091, 0.128, 0.149, and 0.551.

Answer by ikleyn(52858)

.
First, I checked that the given sequence P(0), P(1), P(2), P(3), P(4), P(5), P(6), P(7) of 8 terms monotonically increases 
and that the sum of all 8 (eight) given probabilities in the post is equal to 1 :


    0.001 + 0.006 + 0.022 + 0.052 + 0.091+ 0.128 + 0.149 + 0.551 = 1,


so the given data makes sense.



Second, the answer to the question is the difference


    P(the number of phone calls is from 2 to 6 inclusive) = P(6) - P(0) - P(1) = 


       = the 7-th term of the sequence MINUS the 1-st term and MINUS the 2-nd term  = 

       = 0.149 - 0.001 - 0.006 = 0.142.     ANSWER